# PSN Game: Game-theoretic Prediction and Planning via a Player Selection Network

### v2

@@ 3.3. Player Selection Network (PSN)
In the ideal case, the mask $m^{ij}$ should be binary as defined in (3), which hinders backpropagation and there would be $2^{N-1}$ possible masks $M^{i}$. Although it is possible to train a neural network with categorical outputs via the Gumbel softmax reparameterization Jang et al. (2017), it becomes too expensive when the initial number of players $N$ is large. To overcome such limitations, we set the output masks as a continuous value $m^{ij}\in[0,1]$, and relax the shared cost $c_{s,k}^{i}(\textbf{x}_{k},\textbf{u}_{k};\theta^{i})$} as $\hat{c}_{s,k}^{i}(\textbf{x}_{k},\textbf{u}_{k};\theta^{i})$}, where
@@
This allows us to use a continuous $m^{ij}$ to represent the filtering of shared costs. At runtime, a threshold $m_{\text{th}}$ is introduced to convert the continuous output of PSN to binary masks ($1$ if larger than $m_{\text{th}}$ and 0 otherwise).
@@
$L_{\text{Pred}}$ $=L_{\text{Binary}}+\sigma_{1}L_{\text{Sparsity}}+\sigma_{2}L_{\text{Similarity}}$ (6)
$L_{\text{Plan}}$ $=L_{\text{Binary}}+\sigma_{3}L_{\text{Sparsity}}+\sigma_{4}L_{\text{Cost}}$
where
$L_{\text{Binary}}$ $=\frac{1}{N}\sum_{j=1}^{N}m^{ij}(1-m^{ij}),\quad L_{\text{Sparsity}}=\frac{\|M^{i}\|_{1}}{N},$
$L_{\text{Similarity}}$ $=\sum_{k}\|h(\hat{\textbf{x}}_{k}^{i})-p_{k}^{i}\|_{2},\quad L_{\text{Cost}}=J^{i}(\hat{\textbf{x}}^{i};\textbf{x}^{\neg i}).$
with respective nonnegative weights $\sigma_{1:4}$. Observation $p_{k}^{i}$ is associated with the trajectory of the ego agent $i$ at time $k$, and $\textbf{x}^{\neg i}$ is the other agents’ trajectory if ego agent $i$ were to consider all other agents; in contrast, $\hat{\textbf{x}}^{i}$ is agent $i$’s computed Nash trajectory for the game with relaxed shared cost in (5) (i.e., returned by the aforementioned differentiable game solver). $L_{\text{Binary}}$ encourages mask $m^{ij}$ to converge to either 0 or $1$. $L_{\text{Sparsity}}$ encourages agent $i$ to consider fewer other agents. $L_{\text{Sim}}$ encourages agent $i$ to consider agent subsets which lead to Nash solutions that recover its observed trajectory, and $L_{\text{Cost}}$ encourages agent $i$ to consider appropriate agents to minimize its game cost.
3.4. Goal Inference Network (GIN)
In order to employ the PSN in more complicated scenarios where the game information is incomplete (i.e., the parameter $\theta$ introduced in Section 3.1 is unknown), it is necessary to recover $\theta$ from observations of agents’ states. In this work, we focus on treating $\theta^{i}$ as agent $i$’s 2-D goal position $\hat{p}_{g}^{i}$ (more generally, $\theta^{i}$ can be any parameters, such as weights on basis functions spanning agent preferences, which have been studied in Peters et al. (2021); Qiu and Fridovich-Keil (2024); Liu et al. (2023)). We introduce a data-driven goal inference network $G_{\phi}$, which effectively infers the agent’s goal $\hat{p}_{g}^{i}$ from the (partial) observation of their past trajectories $\{h(\textbf{x}_{k})\}_{k=0}^{K}$. The network is trained from the dataset $\mathcal{D}$, which contains the Nash equilibrium trajectories of all agents from solving a game with ground truth goals, by minimizing the mean goal inference error $L_{\text{Goal}}$ for each sample $d$:
@@ 3.4. Goal Inference Network (GIN)
3.5. Receding Horizon Prediction and Planning
We now integrate the player selection network with a (differentiable) Nash game solver [”Sun”] (2025) in the receding time horizon for prediction and planning tasks.
In prediction tasks, as illustrated in Algorithm 1, the first $K$ steps of the states of all agents are observed, and their goals $\hat{p}_{g}$ are inferred by the GIN. At each step, PSN returns an adaptive mask $M_{k}^{i}\coloneqq(m_{k}^{ij})$, which determines the key players for the interested agent $i$. Then an OLNE strategy $(u_{k|k}^{i*},\dots,u_{k+T-1|k}^{i*})$ is solved from the masked game and the one-step state prediction for agent $i$ is obtained from acting $u_{k|k}^{i*}$, assuming the other agents are still obtained by solving a game with all agents.
In practice, this complexity could be avoided by treating each non-ego player $j\neq i$ as ego and solving a masked game for that player to predict its action. However, for the sake of interpretability, we only employ the PSN for predicting the ego agent’s actions when reporting results in Section 4.
In planning tasks, as illustrated in Algorithm 2, at each time step $k\geq K$, ego agent $i$ observes all agents’ past trajectories of $K$ steps $\textbf{x}_{k-K:k}$, infers their goals $\hat{p}_{g}$ via GIN and obtains an adaptive mask $M_{k}^{i}\coloneqq(m_{k}^{ij})$ from the neural network, which determines the key players. The ego agent then solves for an GOLNE strategy $(u_{k|k}^{i*},\dots,u_{k+T-1|k}^{i*})$ for $\Gamma^{i}(\tilde{\textbf{x}}_{k},\tilde{\textbf{f}};\tilde{\theta},M_{k}^{i})$} and then implements the first control input $u_{k|k}^{i*}$ and updates its state according to (1b). The resulting receding-horizon strategy is denoted as $\textbf{u}_{\text{RH}}^{i*}=(u_{k|k}^{i*},u_{k+1|k+1}^{i*},\dots)$}.
@@ 3.5. Receding Horizon Prediction and Planning
0:  PSN-Prediction (Full, Partial), Goal Inference Network $G_{\phi}$, receding horizon masked game $\Gamma^{i}(\tilde{\textbf{x}}_{k},\tilde{\textbf{f}};\{p_{g}^{i}\}_{i=1}^{N},M_{k}^{i})$, observation interval $K$, prediction interval $T_{\text{Pred}}$.
0:  Agent $i$’s receding horizon trajectory prediction $\hat{\textbf{x}}_{K+1:K+T}^{i}$}.
0:   // Infer agents’ goals via $G_{\phi}$.
1:  ${\color[rgb]{0,0,1}\{\hat{p}_{g}^{i}\}_{i=1}^{N}}\leftarrow G_{\phi}(\{h(\textbf{x}_{k})\}_{k=0}^{K})$ if $\{p_{g}^{i}\}_{i=1}^{N}$ not known, else ${\color[rgb]{0,0,1}\{\hat{p}_{g}^{i}\}_{i=1}^{N}}\leftarrow\{p_{g}^{i}\}_{i=1}^{N}$
2:  for $k=K,\cdots,K+T_{\text{Pred}}-1$ do
2:    // Identify key agents via PSN.
3:   ${\color[rgb]{1,0,0}M_{k}^{i}}\leftarrow\text{PSN-Full}(\hat{\textbf{x}}_{k-K:k})$ or $\text{PSN-Partial}(\hat{\textbf{p}}_{k-K:k})$.
3:    // Solve masked game for agent $i$.
4:   $u_{k|k}^{i*}\leftarrow\text{solving }\Gamma_{k}^{i}(\tilde{\textbf{x}}_{k},\tilde{\textbf{f}};{\color[rgb]{0,0,1}\{\hat{p}_{g}^{i}\}_{i=1}^{N}},{\color[rgb]{1,0,0}M_{k}^{i}})$.
4:    // Solve full game for the other agents.
5:   $u_{k|k}^{-i*}\leftarrow\text{solving }\Gamma_{k}(\textbf{x}_{k},\textbf{f};{\color[rgb]{0,0,1}\{\hat{p}_{g}^{i}\}_{i=1}^{N}})$.
5:    // Update prediction based on dynamics (1b).
6:   $\hat{x}_{k+1}^{i}\leftarrow f^{i}(x_{k}^{i},u_{k|k}^{i*})$.
7:   $\hat{x}_{k+1}^{-i}\leftarrow f^{-i}(x_{k}^{-i},u_{k|k}^{-i*})$.
8:   $k\leftarrow k+1$.
9:  end for
@@
0:  PSN-Planning (Full, Partial), Goal Inference Network $G_{\phi}$, receding horizon masked game $\Gamma^{i}(\tilde{\textbf{x}}_{k},\tilde{\textbf{f}};\{p_{g}^{i}\}_{i=1}^{N},M_{k}^{i})$, observation interval $K$, ego agent $i$’s goal $p_{g}^{i}$.
0:  Agent $i$’s receding horizon strategy $\textbf{u}_{\text{RH}}^{i*}$}.
0:   // Infer the other agents’ goals via $G_{\phi}$.
1:  ${\color[rgb]{0,0,1}\{\hat{p}_{g}^{j}\}_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}}\leftarrow G_{\phi}(\{h(\textbf{x}_{k})\}_{k=0}^{K})$ if $\{p_{g}^{j}\}_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}$ not known,else ${\color[rgb]{0,0,1}\{\hat{p}_{g}^{j}\}_{j=1}^{N}}\leftarrow\{p_{g}^{j}\}_{j=1}^{N}$.
2:  while $k\geq K$ and player $i$ not reaching his goal $p_{g}^{i}$ do
2:    // Identify key agents via PSN.
3:   ${\color[rgb]{1,0,0}M_{k}^{i}}\leftarrow\text{PSN-Full}(\textbf{x}_{k-K:k})$ or $\text{PSN-Partial}(\textbf{p}_{k-K:k})$.
3:    // Solve Masked game for agent $i$.
4:   $u_{k|k}^{i*}\leftarrow\text{solving }\Gamma_{k}^{i}(\tilde{\textbf{x}}_{k},\tilde{\textbf{f}};{\color[rgb]{0,0,1}\{\hat{p}_{g}^{i}\}_{i=1}^{N}}{\color[rgb]{1,0,0}M_{k}^{i}})$.
4:    // Update agent $i$’s state based on dynamics (1b).
5:   $x_{k+1}^{i}\leftarrow f^{i}(x_{k}^{i},u_{k|k}^{i*})$ based on dynamics (1b).
6:   $k\leftarrow k+1$.
7:  end while
4. Experiments
In this section, we address two questions: i) how does the PSN Game perform against baseline selection methods on both prediction and planning tasks; and ii) can PSN Game adapt to challenging settings such as information deficiency (i.e., unknown objectives of non-ego agents), varying number of agents, and real-world pedestrian scenarios? We train and validate PSN Game to investigate the following hypotheses:
Hypothesis 1: In prediction tasks, PSN-Prediction yields more accurate forecasts of the target agents’ future trajectories than baseline selection methods.
Hypothesis 2: In planning tasks, PSN-Planning produces safer and more efficient plans (lower navigation/collision/control costs, smoother and shorter trajectories) than baselines.
Hypothesis 3: PSN readily adapts to incomplete-information games when paired with the goal inference network.
Hypothesis 4: PSN scales to scenarios with more agents than seen in training, without retraining or finetuning.
Hypothesis 5: PSN transfers to real pedestrian trajectories.
4.1. Experiment Setup
4.1.1. Game Structure and PSN Training
In this work, we use the following game setup for training and evaluation: $N$ players are moving towards their goals while avoiding each other. At each time $k$, the $i^{\text{th}}$ player’s state $x_{k}^{i}\in\mathbb{R}^{4}$} encodes its position $p_{k}^{i}=[p_{x,k}^{i},p_{y,k}^{i}]^{\top}$} and velocity $v_{k}^{i}=[v_{x,k}^{i},v_{y,k}^{i}]^{\top}$}, and evolves according to double-integrator dynamics, i.e.,
$x_{k+1}^{i}$ $=\begin{bmatrix}p_{k+1}^{i}\\ v_{k+1}^{i}\\ \end{bmatrix}=\begin{bmatrix}I_{2}&I_{2}\Delta t\\ \textbf{0}&I_{2}\end{bmatrix}\cdot\begin{bmatrix}p_{k}^{i}\\ v_{k}^{i}\\ \end{bmatrix}+\begin{bmatrix}\textbf{0}\\ I_{2}\end{bmatrix}\Delta t\cdot a_{k}^{i}$ (8)
$=\underbrace{Ax_{k}^{i}+B\Delta tu_{k}^{i}}_{f^{i}(x_{k}^{i},u_{k}^{i})},\ k\in[T],\ i\in[N].$
$c_{k}^{i}$ $=c_{p,k}^{i}+c_{s,k}^{i}$ (9a)
$c_{p,k}^{i}$ $=w_{1}^{i}\|p_{k}^{i}-p_{\text{ref},k}^{i}\|_{2}^{2}+w_{2}^{i}\|v_{k}^{i}\|_{2}^{2}+w_{3}^{i}\|u_{k}^{i}\|_{2}^{2}$ (9b)
$c_{s,k}^{i}$ $=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}c_{k}^{ij}=w_{4}^{i}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\exp{\left(-\|p_{k}^{i}-p_{k}^{j}\|_{2}^{2}\right)}.$ (9c)
This form of objective guides the $i^{\text{th}}$ player towards its destination $p_{g}^{i}$ along a goal-oriented reference path $\{p_{\text{ref},k}^{i}\}_{t=1}^{T}$, where
$p_{\text{ref},k}^{i}=(1-\tfrac{k}{T})\,p_{0}^{i}+\tfrac{k}{T}\,p_{g}^{i},$
with minimal velocity ($\|v_{k}^{i}\|_{2}^{2}$) and energy expenditure ($\|u_{k}^{i}\|_{2}^{2}$) while maintaining a safe distance from the others ($\exp{(-\|p_{k}^{i}-p_{k}^{j}\|_{2}^{2})}$).
4.1.2. Tasks and Metrics
We evaluate PSN on prediction and planning tasks. For prediction, we report Average Displacement Error (ADE $\downarrow$), Final Displacement Error (FDE $\downarrow$), and Selection Consistency (Consistency $\uparrow$) between the ego agent’s computed and ground-truth trajectories. For planning, we report Navigation Cost (Nav. Cost $\downarrow$), Collision Cost (Col. Cost $\downarrow$), Control Cost (Ctrl. Cost $\downarrow$), Selection Consistency (Consistency $\uparrow$), Minimum Distance (Distm $\uparrow$), Trajectory Smoothness (trajS $\downarrow$), and Trajectory Length (trajL $\downarrow$). We use $\uparrow$ for larger-is-better and $\downarrow$ for smaller-is-better. Metric implementations are provided in the Supplementary Material.
4.1.3. Baselines
We compare our proposed PSN Game against the following baseline selection methods. In each method, an agent is selected if the following criteria are met:
  1. i.
  2. ii.
    PSN-Rank (ours): the agent has one of the top $|U^{i}|$ largest mask $m^{ij}$.
  3. iii.
  4. iv.
  5. v.
    Gradient Chahine et al. (2023): the agent has one of the top $|U^{i}|$ largest norms of the gradient of the ego agent’s shared cost with respect to the agent’s current control input.
  6. vi.
    Hessian Chahine et al. (2023): the agent has one of the top $|U^{i}|$ largest norms of the Hessian of the ego agent’s shared cost with respect to the agent’s current control input.
  7. vii.
    Cost Evolution Chahine et al. (2023): the agent causes one of the top $|U^{i}|$ largest increases in the ego agent’s shared cost over the last step.
  8. viii.
    BF Chahine et al. (2023): the agent has one of the top $|U^{i}|$ smallest barrier function encoding the safety constraint.
  9. ix.
    CBF Chahine et al. (2023): the agent has one of the top $|U^{i}|$ smallest control barrier function encoding the safety constraint.
4.1.4. Scenarios
To investigate the above hypotheses, we train both PSN-Prediction and PSN-Planning in an idealized setting for 4 and 10 agents with known ground-truth goals. At test time, PSNs are benchmarked against baselines in a Monte-Carlo study over 4 and 10 agent scenarios in prediction and planning tasks with ground truth goals (Hypotheses 1 and 2) and inferred goals (Hypothesis 3). We also evaluate the PSN that was trained for 10 agents in 20 agent scenarios to validate its scalability to a larger number of agents (Hypothesis 4) and finally, to real pedestrian trajectories in the CITR Yang et al. (2019) dataset to assess generalization well outside the training distribution (where human motions are not governed by an explicit test-time game model) (Hypothesis 5). Scenario-specific details are in the Appendix.
4.2. Qualitative Analysis
Advantages of PSN Game: Table 1 summarizes the required information for runtime operation and parameters for different player selection methods, including agents’ latest position (Pos.), latest velocity (Vel.), latest control (Ctrl.), history trajectory (Traj.), and game cost/constraint (Game). Our method offer several advantages: i) PSN-Partial requires only position information, which is more accessible compared to velocities and control inputs; ii) PSN Game reasons about agents’ importance based on trajectory histories, not just instantaneous states, promoting better long-term consistency; iii) while PSN Game does require a model of agents’ cost and constraints during training, that information is not used at runtime; iv) PSN Game requires no environment-specific parameter tuning at runtime. These properties enhance the versatility and general applicability of our framework.
Advantages of Threshold-Based Methods over Ranking-Based Methods: As a threshold-based method, PSN Game provides several advantages over ranking-based methods:
Adaptive selection: Choosing an appropriate threshold (e.g., $m_{\text{th}}=0.5$) is simpler and more robust across scenarios than fixing the number of selected players.
Variable player set: Thresholding allows the number of selected players to adapt naturally to environmental density, unlike ranking-based methods that enforce a fixed number. Trajectory visualization in Fig. 3 illustrates that ranking-based methods may force the ego agent to select irrelevant players to meet a fixed quota, leading to unnecessary computation.
Better flexibility: Threshold-based methods can easily convert to ranking-based selections if needed, but not vice versa due to metric variability across environments.
Table 1. Characteristics of player selection methods

### v3

@@ 3.3 Player Selection Network (PSN)
In the ideal case, the mask $m^{ij}$ should be binary as defined in equation 3, which hinders backpropagation and the combinatorial mask space contains $2^{N-1}$ possible masks $M^{i}$. Although it is possible to train a neural network with categorical outputs via the Gumbel softmax reparameterization (Jang et al., 2017), it becomes expensive when $N$ is large. To overcome such limitations, we set each mask entry to a continuous value $m^{ij}\in[0,1]$, and relax the shared cost $c_{s,k}^{i}(\textbf{x}_{k},\textbf{u}_{k};\theta^{i})$} as $\hat{c}_{s,k}^{i}(\textbf{x}_{k},\textbf{u}_{k};\theta^{i})$}, where
@@
This yields a soft mask that continuously reweights shared interaction costs. At runtime, a threshold $m_{\text{th}}$ converts the continuous PSN output to binary masks ($1$ if larger than $m_{\text{th}}$ and 0 otherwise).
@@
$L_{\text{Pred}}=L_{\text{Binary}}+\sigma_{1}L_{\text{Sparsity}}+\sigma_{2}L_{\text{Similarity}},\quad L_{\text{Plan}}=L_{\text{Binary}}+\sigma_{3}L_{\text{Sparsity}}+\sigma_{4}L_{\text{Cost}}\\$ (6)
$L_{\text{Binary}}=\frac{1}{N}\sum_{j=1}^{N}m^{ij}(1-m^{ij}),\ L_{\text{Sparsity}}=\frac{\|M^{i}\|_{1}}{N},\ L_{\text{Similarity}}=\sum_{k}\|h(\hat{\textbf{x}}_{k}^{i})-p_{k}^{i}\|_{2},\ L_{\text{Cost}}=J^{i}(\hat{\textbf{x}}^{i};\textbf{x}^{\neg i}).$
with respective nonnegative weights $\sigma_{1:4}$. Here, $p_{k}^{i}$ denotes the observed position of ego agent $i$ at time $k$, and $\textbf{x}^{\neg i}$ is the other agents’ trajectory if ego agent $i$ were to consider all other agents; in contrast, $\hat{\textbf{x}}^{i}$ is agent $i$’s computed Nash trajectory for the game with relaxed shared cost in equation 5 (i.e., returned by the aforementioned differentiable game solver). $L_{\text{Binary}}$ encourages mask $m^{ij}$ to converge to either 0 or $1$. $L_{\text{Sparsity}}$ encourages agent $i$ to consider fewer other agents. $L_{\text{Similarity}}$ encourages agent $i$ to consider agent subsets which lead to Nash solutions that recover its observed trajectory, and $L_{\text{Cost}}$ encourages agent $i$ to consider appropriate agents to minimize its game cost.
3.4 Goal Inference Network (GIN)
In order to employ the PSN in more complicated incomplete-information scenarios (i.e., where the parameter $\theta$ introduced in Section 3.1 is unknown), it is necessary to recover $\theta$ from observations of agents’ states. In this work, we focus on treating $\theta^{i}$ as agent $i$’s 2-D goal position $\hat{p}_{g}^{i}$. More generally, $\theta^{i}$ can represent other parameters, such as weights on basis functions spanning agent preferences (Peters et al., 2021; Liu et al., 2023; Qiu and Fridovich-Keil, 2024). We introduce a data-driven goal inference network $G_{\phi}$, which infers an agent’s goal $\hat{p}_{g}^{i}$ from partial observations of that agent’s past trajectory $\{h(\textbf{x}_{k})\}_{k=0}^{K}$. The network is trained from the dataset $\mathcal{D}$, which contains the Nash equilibrium trajectories of all agents from solving a game with ground truth goals, by minimizing the mean goal inference error $L_{\text{Goal}}$ for each sample $d$:
@@ 3.4 Goal Inference Network (GIN)
3.5 Receding Horizon Prediction and Planning
We now integrate the player selection network with a (differentiable) Nash game solver (Sun, 2025) in the receding time horizon for prediction and planning tasks.
In prediction tasks, as illustrated in Algorithm 1, the first $K$ steps of the states of all agents are observed, and their goals $\hat{p}_{g}$ are inferred by the GIN. At each step, PSN returns an adaptive mask $M_{k}^{i}\coloneqq(m_{k}^{ij})$, which determines the key players for the ego agent $i$. Then an OLNE strategy is computed for the masked game, and the one-step prediction for the ego agent is obtained by applying $u_{k|k}^{i*}$. For the non-ego agents, we use actions obtained from the full game.
In practice, this complexity could be avoided by treating each non-ego player $j\neq i$ as ego and solving a masked game for that player to predict its action. To keep evaluation focused, we apply PSN only to the ego agent when reporting prediction results in Section 4.
In planning tasks, as illustrated in Algorithm 2, at each time step $k\geq K$, ego agent $i$ observes all agents’ past trajectories of $K$ steps $\textbf{x}_{k-K:k}$ and obtains an adaptive mask $M_{k}^{i}\coloneqq(m_{k}^{ij})$ from the neural network, which determines the key players. Goals for non-ego agents are inferred once at initialization via GIN in Algorithm 2. The ego agent then solves for a GOLNE strategy $(u_{k|k}^{i*},\dots,u_{k+T-1|k}^{i*})$ for $\Gamma^{i}(\tilde{\textbf{x}}_{k},\tilde{\textbf{f}};\tilde{\theta},M_{k}^{i})$} and then implements the first control input $u_{k|k}^{i*}$ and updates its state according to equation 1b. The resulting receding-horizon strategy is denoted as $\textbf{u}_{\text{RH}}^{i*}=(u_{k|k}^{i*},u_{k+1|k+1}^{i*},\dots)$}.
@@ 3.5 Receding Horizon Prediction and Planning
1:PSN-Prediction (Full, Partial), Goal Inference Network $G_{\phi}$, receding horizon masked game $\Gamma^{i}(\tilde{\mathbf{x}}_{k},\tilde{\mathbf{f}};\{p_{g}^{i}\}_{i=1}^{N},M_{k}^{i})$, observation interval $K$, prediction interval $T_{\mathrm{Pred}}$
2:Agent $i$’s receding horizon trajectory prediction $\hat{\mathbf{x}}_{K+1:K+T_{\mathrm{Pred}}}^{i}$
[…]
4:for $k=K,\ldots,K+T_{\mathrm{Pred}}-1$ do
5:  ${\color[rgb]{1,0,0}M_{k}^{i}}\leftarrow\mathrm{PSN\mbox{-}Full}(\hat{\mathbf{x}}_{k-K:k})$ or $\mathrm{PSN\mbox{-}Partial}(\hat{\mathbf{p}}_{k-K:k})$ $\triangleright$ Identify key agents via PSN
6:  $u_{k|k}^{i*}\leftarrow\text{solution of }\Gamma_{k}^{i}(\tilde{\mathbf{x}}_{k},\tilde{\mathbf{f}};{\color[rgb]{0,0,1}\{\hat{p}_{g}^{i}\}_{i=1}^{N}},{\color[rgb]{1,0,0}M_{k}^{i}})$ $\triangleright$ Solve masked game for agent $i$
7:  $u_{k|k}^{-i*}\leftarrow\text{solution of }\Gamma_{k}(\mathbf{x}_{k},\mathbf{f};{\color[rgb]{0,0,1}\{\hat{p}_{g}^{i}\}_{i=1}^{N}})$ $\triangleright$ Solve full game for the other agents
8:  $\hat{x}_{k+1}^{i}\leftarrow f^{i}(x_{k}^{i},u_{k|k}^{i*})$ $\triangleright$ Update ego-agent prediction via equation 1b
9:  $\hat{x}_{k+1}^{-i}\leftarrow f^{-i}(x_{k}^{-i},u_{k|k}^{-i*})$ $\triangleright$ Update the other agents’ predictions
10:end for
@@
1:PSN-Planning (Full, Partial), Goal Inference Network $G_{\phi}$, receding horizon masked game $\Gamma^{i}(\tilde{\mathbf{x}}_{k},\tilde{\mathbf{f}};\{p_{g}^{i}\}_{i=1}^{N},M_{k}^{i})$, observation interval $K$, ego agent $i$’s goal $p_{g}^{i}$
2:Agent $i$’s receding horizon strategy $\mathbf{u}_{\mathrm{RH}}^{i*}$
[…]
4:while $k\geq K$ and player $i$ has not reached its goal $p_{g}^{i}$ do
5:  ${\color[rgb]{1,0,0}M_{k}^{i}}\leftarrow\mathrm{PSN\mbox{-}Full}(\mathbf{x}_{k-K:k})$ or $\mathrm{PSN\mbox{-}Partial}(\mathbf{p}_{k-K:k})$ $\triangleright$ Identify key agents via PSN
6:  $u_{k|k}^{i*}\leftarrow\text{solution of }\Gamma_{k}^{i}(\tilde{\mathbf{x}}_{k},\tilde{\mathbf{f}};\{p_{g}^{i},{\color[rgb]{0,0,1}\hat{p}_{g}^{-i}}\},{\color[rgb]{1,0,0}M_{k}^{i}})$ $\triangleright$ Solve masked game for agent $i$
7:  $x_{k+1}^{i}\leftarrow f^{i}(x_{k}^{i},u_{k|k}^{i*})$ $\triangleright$ Update agent $i$’s state via equation 1b
8:  $k\leftarrow k+1$
9:end while
4 Experiments
Table 1: Characteristics of player selection methods